### v1

# Modeling Light Signals Using Data from the First Pulsed Neutron Source Program at the DUNE Vertical Drift ColdBox Test Facility at CERN Neutrino Platform

Abstract
In this paper, we present a first quantitative test of detected light signals produced in a pulsed neutron source run in a small vertical drift LArTPC at the CERN neutrino platform ColdBox test facility. The ColdBox cryostat, detectors, neutron sources, and particle interactions are modeled and simulated using Fluka. A good agreement is found in the detected number of photoelectrons, with values below 650 photoelectrons in both data and simulation, for all four X-ARAPUCA photodetectors on the cathode in the LArTPC. A time constant is also fitted from the neutron-beam-off light signal spectrum and found consistent between data and MC. Several important systematic effects are discussed and serve as guides for future runs at larger LArTPCs.
1 Motivation
The Deep Underground Neutrino Experiment (DUNE) [DUNE_Physics_TDR] is a next-generation long-baseline neutrino oscillation experiment in the US. It will have four far detector (FD) modules, each holding 17 kilotons of liquid argon (LAr). These modules are located 1500 meters underground at the Sanford Underground Research Facility (SURF), South Dakota, and about 1300 kilometers from the near detector complex at Fermilab, Chicago. The vertical drift (VD) FD module [DUNE_VD_TDR] will be the first of the four FD modules to be installed at SURF and is under construction. The VD LAr time projection chamber (TPC) features X-ARAPUCA (XA) photon detectors [Arapuca_2016, XA_2018] installed on the cathode plane and the cryostat membrane. The VD photon detection system (PDS) provides accurate timing of an event (so-called ), background rejection, and calorimetric measurement capabilities. The R&D and validation of the VD LArTPC detector technologies have been enabled by the Neutrino Platform at CERN through its ColdBox and ProtoDUNE programs, which have been an extraordinary success. More details on ProtoDUNE, ColdBox, and future planned tests can be found in ref. [ProtoDUNE3].
MeV energy scale calibration of the PDS is crucial for DUNE’s low energy physics program. One promising candidate calibration technique utilizes the neutron capture on 40Ar. Once the neutron is captured, the gamma cascade from the excited 41Ar has a total energy release of 6.1 MeV, which may serve as a standard candle for PDS calorimetry calibration. By tagging neutron captures, it is possible to derive a reliable light yield map for the energy measurement of MeV neutrino events. This paper presents the first study of light signals produced by neutron interactions in LAr, including their capture on 40Ar, using a pulsed neutron source (PNS) [DUNE_SP_TDR].
2 VD ColdBox Test Facility at CERN Neutrino Platform
The PNS program is performed at the CERN VD ColdBox (CB) test facility. The CB cryostat has an internal volume of 3.89 m (y) × 3.91 m (z) × 1 m (x) and is filled with LAr to a height of 70 cm during normal operation. The drift direction is in +x. A high voltage of -10 kV is applied to a cathode of size 3.37 m x 2.98 m in y-z plane, establishing a uniform electric field of 454 V/cm in the drift region. The total drift time is 140 s for a 21.5 cm drift distance. Four 60 cm 60 cm-sized XA photon detectors are assembled at CIEMAT and NIU. They are installed on the cathode surface facing the anode charge readout plane. The front-end cold readout electronics of these four XA detectors are powered using the Power-over-Fiber (PoF) technology [DUNE_PoF], which enables safe and noise-free detector operation on the high-voltage cathode surface at the cryogenic temperature. The photodetector layout in the CB is shown in Fig. 1 and implemented in Fluka geometry as shown in Fig. 3. A combination of different types of SiPMs is used in the four XA photon detectors. C1 is equipped with FBK SiPMs featuring ‘Metal-in-Trench’ technology, while C2, C3, and C4 are equipped with HPK SiPMs. All SiPMs are operated at 5 V above their breakdown voltage. Two more XAs are also installed on one side of the membrane. However, they are not considered in this study.
3 Pulsed Neutron Source
The PNS used in the run is a commercial Thermo Scientific MP 320 Deuterium-Deuterium generator (DDG). It is deployed on one side of the CB at the mid-height of the drift region. Custom polyethylene blocks are used as neutron shielding for radiation safety. Lead blocks and borated polyethylene are placed in front of the DDG to absorb gammas produced from the polyethylene. The DDG produces 1 million mono-energetic 2.5 MeV neutrons per second. However, in this run at the CB, the DDG is operated in a special mode (Fig. 2), where five bursts of neutrons are generated every 12.5 ms (80 Hz). In this special mode, the actual emission intensity is unknown. Each burst lasts 60 s followed by a 20 s long downtime. Data acquisition rate is 4 Hz. The produced neutrons penetrate the CB wall and the non-instrumented LAr region before reaching the bulk LAr with active charge readout. Neutron interaction products reaching the dead LAr region can produce scintillation light and can be detected by XAs.
4 Samples
4.1 Data
PNS data were acquired with both the TPC and PDS read out using an external trigger provided by the DDG synchronous with the neutron beam. For the analysis, only runs with substantial statistics were considered, resulting in a total of over 160,000 PNS triggers, each recorded with a 1-ms readout window. More than 250,000 cosmic ray triggers were collected during the same period, using a 4-ms readout window. These cosmic datasets were used for calibration and background estimation. Test runs with limited statistics or used for debugging purposes were excluded from the study. These are summarized in Tab. 1.
| Type | Triggers | Window (ms) |
| PNS | 162908 | 1 |
| Cosmic | 253734 | 4 |
4.2 Monte Carlo
The simulation of the PNS at the CB is implemented in Fluka2024.1 [Fluka] as shown in Fig. 3, taking advantage of Fluka’s good modeling of neutron transport and nuclear deexcitation. In the simulation, each event corresponds to a single neutron from the DDG, and all neutrons are emitted at time zero. Optical photons are produced and transported both in the active readout LAr volume and the non-active regions. The adopted scintillation light yield is 2.55104 photons/MeV, as estimated by LArQL[LArQL] for a Minimum Ionizing particle (MIP) in an electric field of 454 V/cm. The same scintillation photon yield is used for both the active and inactive LAr regions. LArQL has been validated on data at 500 V/cm, where the expected light yield is 2.4104 photons/MeV. The fast and slow time constants (6 ns and 1.5 s) of Ar scintillation light are used. A Rayleigh length of 90 cm is adopted as well as an infinite absorption length [Fluka_validation]. A photon detection efficiency (PDE) of 3% is used for all XAs. To decrease CPU time, only 1/10 of the optical photons are tracked. These scale factors are taken into account later in the analysis of the MC sample.
Typical physics processes relevant for neutrons from the PNS include neutron inelastic interactions, neutron elastic interactions, and neutron captures. Each process could leave energy deposits in LAr and generate scintillation light. Events where there is either energy deposition in active LAr, or one neutron capture in active LAr, or scintillation photons arriving at the XAs from neutron interactions on any inactive materials, are recorded. For each recorded event, interactions occurring in all regions, the associated energy depositions, and optical photons are all stored. The recorded events correspond to approximately 3.2% of generated neutrons. Events with one neutron capture are 0.04% of the generated neutrons.
Fluka performs a detailed simulation of neutron transport and interaction, including the description of gamma ray cascades according to the available nuclear data [Fluka]. Fig. 4 shows the spectrum following neutron capture in Ar in the Fluka simulation. The most intense emissions, at 4.7 MeV, 1.2 MeV, and 167 keV, are clearly visible. Due to the small active volume of the detector, neutron interactions in the surrounding inactive materials have to be carefully modeled. Fig. 4 also shows the detailed simulated spectrum from non-Ar materials, showing, for instance, the 2.2 MeV line from capture on hydrogen, the 7.6 MeV line from capture on Fe, and the 876 keV line from inelastic scattering on Fe.
A good representation of what happens outside the readout volume is particularly important in the present setup, where the readout volume is relatively small. Table 2 shows the fraction of events associated with the interaction region. Indeed, most of the signals collected by the most exposed XA (C4) originate from neutron interactions occurring outside the active LAr region.
| Region | light and charge | light only |
| Active LAr | 0.29 | 0.29 |
| Non-active LAr | 0.26 | 0.26 |
| PNS shielding polyethylene | 0.19 | 0.19 |
| Cryostat outer wall | 0.16 | 0.17 |
| PNS tube | 0.03 | 0.03 |
| Cryostat inner wall | 0.025 | 0.025 |
| Other | 0.04 | 0.03 |
5 Calibration and Data Selection
In this section, we present calibration procedures for the amplitude of light signals on each of the four XAs on the cathode.
5.1 ADC to Photoelectron Calibration in data
The digitized counts from analog-to-digital converters (ADC) on XA readout are converted to photoelectrons (PE) via a calibration constant. The calibration constant is dependent on run and channel for a specific XA. They are obtained by finding the average ADC counts for many single PE signals on each XA channel.
5.2 Relative PDE Calibration in data
Different XA configurations, SiPMs, and dichroic filters from different manufacturers could cause variations in the PDE among XA modules. The XA modules in these ColdBox runs are prototypes and don’t represent the final detector design for DUNE. Variations in the PDE of all XAs are calibrated by studying cosmic tracks passing through the center of the CB at close to zero zenith angle (within 5 cm between top and bottom crossing points). For these cosmic tracks, we expect a similar number of photons to arrive at the surface of each of the four XAs on the cathode. Therefore, the detected PEs on each XA provide a calibration of the relative PDE among the four XAs. Tab. 3 shows the average detected PEs on the four XAs from these cosmic tracks. The relative PDE translates to a scale factor in PE, later applied to data on each XA when comparing to MC. The XA module C3 is used as the reference detector, where the PE scale factor is set to one. The other XAs have a scale factor derived from the ratio of their total detected PE to that on C3.
| XA Module | ch0 detected PE | ch1 detected PE | Tot. detected PE | PE scale factor |
| C1 | 125.5 | 130.8 | 256.3 | 1.85 |
| C2 | 68.2 | 62.97 | 131.2 | 0.947 |
| C3 | 69.14 | 69.4 | 138.5 | 1.0 (reference XA) |
| C4 | 57.42 | 48.4 | 105.8 | 0.764 |
5.3 Photon to Peak ADC Calibration in MC
As described in Sec. 5.1, the number of photoelectrons in data is derived from the peak ADC after applying an ADC to PE scale factor. The Fluka simulation provides the number of photons generated in a typical neutron event. These photons are summed and convoluted with the single photoelectron template obtained from LED calibration runs. The peak ADC is then used to calculate the number of photoelectrons, as consistently performed in data. This factor is found to be 0.3, close to the fast component of LAr scintillation light.
5.4 Data Selection
We use the Lardon framework [lardon] for this CB data analysis. A peak finding algorithm is used to find photon signals larger than 15 PE. Due to the different frameworks used in data (Lardon) and simulation (Fluka), we choose to apply a minimal selection for an unbiased data-MC comparison. Future work will seek to unify the framework so that consistent and more complex event selections can be applied to both.
For two channels on each XA, the signals whose peak times are within 80 ns (5 PD time ticks) are used to calculate the total detected signal for a specific activity. This time coincidence requirement is derived from the peak time difference between all possible pairs of PD signals from two channels on the XA from a cosmic run. For each channel of the XA on the cathode, the ADC-PE calibration constants are used to calculate the PE. The total detected PE on an XA is then obtained by summing up the PE from two channels.
Because of geometric acceptance, activities closer to the XA detector produce a larger number of PE, while those at the center of the CB generate weaker signals. At a small number of PE, the PE count is subject to fluctuations in the waveform baseline, dark counts, and background light noise. As a result, we restrict the region of interest to 100 PE and above for each XA module to avoid analyzing ambiguous low PE pulses. For each channel of XA, we also apply a max ADC cut of 14,000 to remove saturated signals.
6 Result
6.1 Light Signal Amplitude
A full comparison of data from all PNS runs in Tab. 1 to the Fluka simulation is shown in Fig. 5. The cosmic background is modeled with a data-driven method. The data from all cosmic runs are scaled based on the total collected triggers relative to all PNS runs in Tab. 1. Only the first millisecond of data is used for all cosmic runs to be consistent with the PNS trigger window. A scale factor of 0.642 is determined.
A scale factor of 0.3 is applied to the total photons in each simulated neutron event to be consistent with the photoelectron calculation method in data as described in Sec. 5.3. As illustrated in Tab. 3, the C3 module offers the best data quality where both channels have the closest response to cosmic rays. We scale the Fluka simulation to the rest of the data observed on the C3 module (i.e., excluding the cosmic prediction), making perfect agreement. The scale factor is determined to be 5.0511 and applied to MC events for all XAs modules.
The data-to-MC comparison after subtracting cosmic backgrounds from the PNS data is shown in Fig. 6. An overall good agreement between data and MC is observed for all XA modules on the cathode up to about 650 PE. Above 650 PE, an excess of data is observed on all XA modules that is not accounted for by the simulation. This could originate from several possible sources, including but not limited to poorly modeled cosmic background, unmodeled neutron-induced interactions within the XA modules, and the accidental coincidence of cosmic ray activity with neutron interactions in LAr. These processes are not considered in the simulation. Further study on these possible sources is presented in Sec. 6.3.
The small disagreement below 650 PE could come from minor displacement of detector positions or the position of the neutron source in the simulation, as part of the systematic effects discussed in Sec. 7. For modules that are closer to the PNS (C4 being the closest), a larger neutron-induced background from interactions on materials outside active LAr is observed. The neutron inelastic scattering background in the active LAr region is small on all modules, as we expect most inelastic interactions to happen earlier in the CB wall and the passive LAr region. For the C1 module, the data stops around 1100 PE due to the applied relative PDE calibration factor of 1.85 as described in Sec. 5.2. The total number of events for each component in Fig. 5 is listed in Tab. 4.
| XA | Cosmics | Inactive LAr (MC) | Active LAr Inelastic (MC) | Active LAr Capture (MC) | Total Predicted (MC) | Data | Data - Cosmics | <PE> |
| C1 | 111983 | 36939 | 2101 | 24988 | 64028 | 161979 | 49996 | 286 |
| C2 | 94119 | 7860 | 248 | 4970 | 13077 | 112471 | 18352 | 401 |
| C3 | 92630 | 8718 | 298 | 6864 | 15881 | 108511 | 15881 | 366 |
| C4 | 83828 | 143290 | 9723 | 46788 | 199802 | 205299 | 121471 | 343 |
6.2 Light Signal Time Constant
Fig. 7 shows an observed beam bunch structure in data on the C4 module closest to the neutron source. A similar beam time profile is observed in all PNS runs, with a slight shift of the observed number of bunches due to data acquisition time. An exponential function is fitted to the decay part of the spectrum between 12000 and 60000 time ticks (i.e., 192 - 960 s), and a decay time constant of 283 12 s is obtained.
To further understand the processes that contribute to the exponential decay tail in the timing spectrum, we first reproduce the timing structure observed in the data. We assign a random time to the original neutron in the Fluka simulation, assuming a time structure with 5 bunches, each one 60 s separated by 20 s, of which only the last one and a half are in the trigger window. Fig. 7 right plot shows the time distribution of optical photons from all simulated neutron events. The characteristic decay time constant is fit to be 254.5 68 s in the Fluka simulation from the decay spectrum between 100 and 900 s, consistent with that obtained in the fit to data within one standard deviation. Fluka simulation also suggests the decay time spectrum receives primary contribution from the neutron interactions outside the active LAr volume.
6.3 Discussion on Excess at High PE in Data
Here we discuss several possible sources that could produce the high PE excess observed in the data in Fig. 6.
Pile up light signals. Pile-up (PU) light signals occur when two events happen close in time. In this analysis, the signal amplitude could be overestimated when the signal overlaps with the long tail of an earlier pulse. This could happen, for example, when a cosmic muon precedes a neutron-induced signal. We performed a data-driven estimation by tagging and separating the PU light pulses. Because of the intrinsic slow scintillation light component (1.5 s) in LAr, a 10 s-long moving window is used to identify PU light pulses that are contained in this timing window for each channel on the XA. All pulses except the earliest pulse in this moving time window are tagged as PU pulses. The same data selection in Sec. 5.4 is applied. Overall, we found PU pulses constitute about 20% of all light signals. An overlay of the signal amplitude of the PU tagged pulses and the rest pulses on the C4 XA module is shown in Fig. 8. The similar shapes for PU pulses and the rest indicate PU is not causing the excess observed in high PE.
Cosmic background. The excess at high PE region could originate from not well modeled cosmic flux from cosmic runs. To test this hypothesis, we performed a fit to the timing distribution separately for events with number of PE below and above 1200 in Fig. 6. For events with PE 1200, the timing distribution would be flat as one would expect for cosmic background. Instead, a decay time profile similar to the events with PE 1200 is observed (Fig. 9). This provides strong evidence that the excess at high PE is from neutron-related signals not modeled in the simulation.
Cherenkov light from charged particles crossing the XA detector. Charged particles entering the plastic layer of the XA detector produce Cherenkov light. The collection and detection efficiency of this light is completely unknown. A test simulation has been performed to evaluate the relevance of this effect. Even assuming a perfect detection efficiency, those events would contribute to about 2% of the signal above 100 PE, and the events producing more than 1200 PE would be about 0.2% of the integrated signal, much smaller than the observed high PE tail.
Electric field outside active LAr. The uncertainty of the electric field in the buffer LAr region is one of the systematics discussed in Sec. 7. In the nominal case, the simulation takes the same 454 V/cm for the buffer LAr region as the active LAr region. In Fig. 10, the change to zero field in the buffer LAr region introduces a high PE tail extending to 2000 PE. This effect could partially explain the high PE excess in the data, as the actual E field in the buffer region is typically smaller than 454 V/cm and is often characterized by complicated field lines that are difficult to model in simulation.
XA Absolute PDE. In the analysis (Sec. 5.2), a relative PDE calibration was performed to the 3% absolute PDE used in the simulation. However, the actual XA absolute PDE can vary by up to 50% when SiPMs are biased at higher overvoltages with a non-negligible cross-talk probability. The absolute PDE is also affected by the detailed XA configuration. A smaller absolute PDE in simulation would underestimate the number of PE and could partially explain the high PE excess in data. This is further discussed in Sec. 7 as a systematic effect. The precise measurement of absolute PDE for XA modules requires a dedicated laboratory setup and analysis. More details can be found in ref. [VD_XA_PDE].
7 Systematic Errors
Several important systematic effects are studied using the Fluka simulation. The uncertainty of the electric field in the buffer LAr region affects the scintillation photon yield. A simulation with zero electric field ("0 field") in this volume is evaluated, where the light yield reaches a maximum compared to the nominal electric field of 454 V/cm. The uncertainty of the initial production of scintillation light in buffer LAr and propagation to the active LAr readout volume is evaluated by setting the photon production in the buffer LAr volume to zero ("act only"). The uncertainty of the PNS position relative to the active volume inside the CB is modeled by shifting the PNS in the x (drift) and y (horizontal) directions by 5 cm ("dispx" and "dispy"). The uncertainty on the photon detector efficiency is studied by increasing it to 4.44% ("heff"), a value that corresponds to the product of efficiency and cross-talk probability for one of the device configurations described in ref. [VD_XA_PDE].
The resulting distribution of the total number of photoelectrons detected by each XA on the cathode under these scenarios is shown in Figure 10. The largest systematic effect is the electric field in the buffer LAr region, which produces almost a factor of two more total PE on the C4 XA module, the closest module to the source, as also shown in the zoom-in plot in Fig. 11. A high detection efficiency obviously increases the high PE tail of the spectra, and the average number of detected PE, especially for the most exposed detectors.
The integrated and average of all photons normalized to the standard simulation condition for each systematic effect is shown in Fig. 12. Overall, the displacement of the neutron generator gives small variations in the arriving optical photons. On the other hand, the field conditions in the non-readout LAr volume have a large impact on XA modules nearest to the source, especially for small signals (left plot in Fig. 12). At a higher number of photons, the electric field impact is smaller, and the modeling of interactions outside active LAr becomes important (right plot in Fig. 12 ).
The peak ADC amplitude to photoelectron calibration constant used in the data discussed in Sec. 5.1 affects the observed amount of light on each XA. The systematic shift of this constant up and down by one ADC unit is applied to the calculation in data, and the resulting counts are presented in Tab. 5. Overall, we observe that the non-cosmic component in data has a max shift of 8.3% to the nominal scenario from C4.
| XA | Cosmics | Inactive LAr (MC) | Active LAr Inelastic (MC) | Active LAr Capture (MC) | Total Predicted (MC) | Data | Data - Cosmics |
| C1 | 106951 | 35511 | 2020 | 24022 | 61552 | 153597 | 46646 |
| C2 | 91296 | 7556 | 238 | 4778 | 12572 | 108528 | 17232 |
| C3 | 88927 | 8381 | 287 | 6599 | 15267 | 104194 | 15267 |
| C4 | 79768 | 137749 | 9347 | 44979 | 192076 | 192214 | 112446 |
| XA | Cosmics | Inactive LAr (MC) | Active LAr Inelastic (MC) | Active LAr Capture (MC) | Total Predicted (MC) | Data | Data - Cosmics |
| C1 | 117967 | 38898 | 2213 | 26313 | 67424 | 171432 | 53465 |
| C2 | 96981 | 8276 | 261 | 5234 | 13771 | 116530 | 19549 |
| C3 | 95721 | 9181 | 314 | 7229 | 16723 | 112444 | 16723 |
| C4 | 88300 | 150889 | 10239 | 49270 | 210398 | 219858 | 131558 |
8 Summary and Outlook
A pulsed neutron source is deployed for the first time at the VD CB facility at the CERN neutrino platform to study MeV neutron interactions with LAr. In this paper, an analysis of light signals is presented. A good modeling of the light signals up to 650 photoelectrons per XA module from the PNS run is demonstrated. A good agreement is found in the fitted decay time constant from neutron interactions after the neutron beam stops.
The modeling of signals from neutron interactions will benefit the DUNE low energy physics program in many ways. MeV-scale physics measurements using scintillation light in neutrino LArTPC are scarce. Efficient tagging of the neutron captures could improve energy resolution [APEX_lowE, MeVNueArCC]. Meanwhile, understanding the neutron capture signals in charge and light will help reject cavern neutron backgrounds. Developing such a neutron capture tagging technique at small LArTPC prototypes before DUNE FD operation is crucial for advancing the DUNE physics program [DUNE_lowE_1, DUNE_lowE_2, DUNE_lowE_3].
Several constraints exist at the CB facility for the physics analysis. The limited drift distance and active volume make it inefficient to contain and tag neutron capture events. However, these constraints are foreseen to be gone in future runs at larger prototypes such as ProtoDUNE-VD. The developed analysis method in this study will be applied to future PNS physics runs in larger LArTPCs. This study opens new possibilities to achieve tagging of capture events for light calorimetry calibration and measurement of the characteristic time of neutron capture on argon.
Acknowledgments
The vertical drift ColdBox test facility was constructed and operated on the CERN Neutrino Platform. We thank the CERN management for providing the infrastructure for this experiment and gratefully acknowledge the support of CERN. We thank LANL for the availability of the DD generator for this physics program.